# Quantum-optical theory of the few femtosecond nonlinear optical response of Drude metals with a non-parabolic conduction band

### v1

@@ Abstract
We develop an energy-space density matrix framework to investigate the interaction of extremely short optical pulses (ESPs) with transparent conducting oxides (TCOs). This approach captures not only electron populations, material polarization, and the permittivity, but also the the quantum coherences between states. Compared to traditional momentum-space models, the energy-space formulation offers substantial computational simplification while retaining accuracy. Building on but going beyond the scope of Ref. [1], we focus on dynamical features previously unexplored. Our formulation reveals clear signatures of quantum coherence in the net absorption dynamics and highlights the emergence of strong excited-state absorption under intense excitation. It also clarifies that spontaneous emission can be neglected in this regime. Furthermore, we investigate the influence of pump pulse intensity on the local field’s duration, spectral broadening and shift, and phase induced by carrier dynamics, highlighting the absorptive nature of the nonlinear response. Our results provide a unified framework for understanding nonlinear light-matter interaction in dispersive, low-density electron systems driven far from equilibrium by intense broadband excitation.
@@ I Introduction
Initial results on this problem were published in [1], which focused on the characterization of the strength of the nonlinear response (in terms of the electron excitation dynamics, material polarization and permittivity). Instead, here, we focus on the formulational aspects of the problem, as well as on additional aspects of the dynamics that were not considered in [1]. Specifically, in Section II, we provide a detailed derivation of the DM formulation that underlies [1], including the transition dipole matrix element calculation in momentum space, and its transformation to the far more computationally-economic energy-space. Then, in Section III, we perform a thorough comparison of the strengths of the various terms appearing in the DM equations and analyze the dynamics of the coherence between the various electron states as well as the resulting repeated frequency doubling dynamics. Then, we identify the signature of the coherence in the (net) energy-integrated absorption dynamics and the emergence of significant excited state absorption within its time-integrated energy spectra; the latter is responsible for the super-linear nature of the absorption, hence, counteracting the effects of population inversion. We also justify the neglect of spontaneous emission (see Appendix A).
@@ II.1 A derivation of a momentum-conserving energy-space density matrix formulation
Here, $\hat{\mathcal{H}}_{\textrm{e,0}}$ is the single-particle approximation for the Hamiltonian of the electrons in the lattice periodic potential, $\hat{\mathcal{H}}_{\textrm{ph,0}}$ is the unperturbed Hamiltonian of the phonon subsystem and $\mathcal{H}_{\textrm{photon-e}}$ is the electric dipole interaction Hamiltonian
$\hat{\mathcal{H}}_{\textrm{photon-e}}=-\hat{d}\cdot E(t),$ (2)
with $\hat{d}$ being the electric transition dipole operator and $E(t)$ being the local electric field, calculated self-consistently with the density matrix elements, see Section II.3. In that regard, we neglect electron- and phonon-assisted absorption events as well as interband transitions in favor of Landau-damping-based absorption (see discussion of Ref. [26] on the analysis of Ref. [27]). Additionally, we neglect the contribution of spontaneous emission (photoluminescence), see justification in Appendix A. $\hat{\mathcal{H}}_{\textrm{e-e}}$ is the electron-electron interaction via the Coulomb potential, $\hat{\mathcal{H}}_{\textrm{e-ph}}$ is the electron-phonon interaction via the deformation potential and $\hat{\mathcal{H}}_{\textrm{e-imp}}$ is the electron-impurity interaction.
@@
$\dot{\rho}(t)=-\dfrac{i}{\hbar}\left[\hat{\mathcal{H}}_{0}+\hat{\mathcal{H}}_{\textrm{photon-e}},\rho\right]+\mathcal{R}(\rho),$ (4)
@@
$\begin{aligned} \dot{f}_{{\bf k}}&=\mathcal{R}_{{\bf k}{\bf k}}(\rho(t))+\dfrac{i}{\hbar}E(t)\sum_{{\bf k}^{\prime}}\left(d_{{\bf k}^{\prime}{\bf k}}\rho_{{\bf k}{\bf k}^{\prime}}-d_{{\bf k}{\bf k}^{\prime}}\rho_{{\bf k}^{\prime}{\bf k}}\right),\end{aligned}$ (5a)
$\begin{aligned} \dot{\rho}_{{\bf k}{\bf k}^{\prime}}=\mathcal{R}_{{\bf k}{\bf k}^{\prime}}(\rho(t))-\dfrac{i}{\hbar}\left(\mathcal{E}_{{\bf k}}-\mathcal{E}_{{\bf k}^{\prime}}\right)\rho_{{\bf k}{\bf k}^{\prime}}-\dfrac{i}{\hbar}E(t)d_{{\bf k}{\bf k}^{\prime}}\left(f_{{\bf k}}(t)-f_{{\bf k}^{\prime}}(t)\right)\\ +\dfrac{i}{\hbar}E(t)\sum_{{\bf k}^{\prime\prime}\neq{\bf k},{\bf k}^{\prime}}\left(d_{{\bf k}{\bf k}^{\prime\prime}}\rho_{{\bf k}^{\prime\prime}{\bf k}^{\prime}}-\rho_{{\bf k}{\bf k}^{\prime\prime}}d_{{\bf k}^{\prime\prime}{\bf k}^{\prime}}\right)\end{aligned}$ $\text{for }{\bf k}\neq{\bf k}^{\prime},$ (5b)
where $\omega_{{\bf k}{\bf k}^{\prime}}\equiv\left(\mathcal{E}_{{\bf k}}-\mathcal{E}_{{\bf k}^{\prime}}\right)/\hbar$ with $\mathcal{E}_{{\bf k}}$ and $\mathcal{E}_{{\bf k}^{\prime}}$ being the renormalized energy of individual electrons and $d_{{\bf k}{\bf k}^{\prime}}$ are the matrix elements of the electric dipole transition operator.
We determine the momentum space dipole transition matrix elements $d_{{\bf k}{\bf k}^{\prime}}$ from the matrix elements of the momentum operator $p_{{\bf k}{\bf k}^{\prime}}$ using the relation
$(d_{x,y,z})_{{\bf k}{\bf k}^{\prime}}=-\dfrac{2ie\hbar}{(p^{2})_{{\bf k}{\bf k}}-(p^{2})_{{\bf k}^{\prime}{\bf k}^{\prime}}}(p_{x,y,z})_{{\bf k}{\bf k}^{\prime}},$ (6)
where $e$ is the electron charge. Eq. (6) is derived from the commutation relation between the position operator and the operator $\hat{{\bf p}}^{2}$ (see [35, Ch. 5]), while the matrix elements of the momentum operator $p_{{\bf k}{\bf k}^{\prime}}$ are determined by assuming an infinitely deep 1D potential well representing the sample geometry, as in popular earlier works [36, 37].
@@
Unfortunately, as shown in Ref. [41], using Eq. (7b) in the equations of motion Eq. (5) leads to a violation of the charge density continuity equation, as well as to an incorrect conductivity in the static limit [42, 43]. These inconsistencies can be resolved by constructing the relaxation operator so that it satisfies $\textrm{Tr}\left(\hat{{\bf d}}\hat{\mathcal{R}}\right)=0$ (see details in Ref. [42, 41]). Following the procedure shown in Refs. [42, 41], a relatively simple relaxation operator preserving the continuity equation can be obtained by replacing Eq. (7b) with
@@
where $h_{{\bf k}{\bf k}^{\prime}}$ is a unitless phase factor such that $h_{{\bf k}{\bf k}^{\prime}}d_{{\bf k}^{\prime}{\bf k}}=d_{{\bf k}{\bf k}^{\prime}}$  [^2] In general, $d_{{\bf k}^{\prime}{\bf k}}$ and $d_{{\bf k}{\bf k}^{\prime}}$ differ by a phase factor. For our choice of a 1D symmetric potential, we have $d_{{\bf k}^{\prime}{\bf k}}=d_{{\bf k}{\bf k}^{\prime}}$..
@@
$\dot{\rho}_{{\bf k}{\bf k}^{\prime}}=\mathcal{R}_{{\bf k}{\bf k}^{\prime}}(\rho(t))-\dfrac{i}{\hbar}\left(\mathcal{E}_{{\bf k}}-\mathcal{E}_{{\bf k}^{\prime}}\right)\rho_{{\bf k}{\bf k}^{\prime}}-\dfrac{i}{\hbar}E(t)d_{{\bf k}{\bf k}^{\prime}}\left(f_{{\bf k}}(t)-f_{{\bf k}^{\prime}}(t)\right).$ (9)
@@
$\begin{aligned} \dfrac{\partial f_{{\bf k}}(t)}{\partial t}=\left(\dfrac{\partial f_{{\bf k}}}{\partial t}\right)_{\textrm{e-e}}+\left(\dfrac{\partial f_{{\bf k}}}{\partial t}\right)_{\textrm{e-ph}}+\left(\dfrac{\partial f_{{\bf k}}}{\partial t}\right)_{\textrm{e-imp}}-\dfrac{i}{\hbar}E(t)\sum_{{\bf k}^{\prime}}\left[d_{{\bf k}^{\prime}{\bf k}}\rho_{{\bf k}{\bf k}^{\prime}}(t)-d_{{\bf k}{\bf k}^{\prime}}\rho_{{\bf k}^{\prime}{\bf k}}(t)\right]\end{aligned}$ (10a)
$\begin{aligned} \dfrac{\partial\rho_{{\bf k}{\bf k}^{\prime}}(t)}{\partial t}=-\eta_{{\bf k}{\bf k}^{\prime}}\left(\rho_{{\bf k}{\bf k}^{\prime}}-h_{{\bf k}{\bf k}^{\prime}}\rho_{{\bf k}^{\prime}{\bf k}}\right)-\dfrac{i}{\hbar}\left(\mathcal{E}_{{\bf k}}-\mathcal{E}_{{\bf k}^{\prime}}\right)\rho_{{\bf k}{\bf k}^{\prime}}-\dfrac{i}{\hbar}E(t)d_{{\bf k}{\bf k}^{\prime}}\left[f_{{\bf k}}(t)-f_{{\bf k}^{\prime}}(t)\right],\end{aligned}$ (10b)
The polarization density can be obtained by taking the trace of the matrix multiplication of the density matrix and the dipole transition matrix, namely,
$P(t)=\dfrac{2}{V}\sum_{{\bf k},{\bf k}^{\prime}}d_{{\bf k}^{\prime}{\bf k}}\rho_{{\bf k}{\bf k}^{\prime}}(t).$ (11)
@@ II.2 From $k$ to $\mathcal{E}$ space formulation
and multiply Eq. (10b) by $d_{{\bf k}^{\prime}{\bf k}}$ and integrate over the solid angles of $\bf k$} and of $\bf k$$^{\prime}$ (see details in Ref. [28]). The photon-e interaction term, the off-diagonal density matrix equation of motion Eq. (10b), and the polarization density Eq. (11) thus, respectively, become
@@
with $\overline{d_{\mathcal{E}^{\prime}\mathcal{E}}\rho_{\mathcal{E}\mathcal{E}^{\prime}}}\equiv\dfrac{1}{(4\pi)^{2}}\displaystyle\int d\Omega_{{\bf k}}d\Omega_{{\bf k}^{\prime}}d_{{\bf k}^{\prime}{\bf k}}\rho_{{\bf k}{\bf k}^{\prime}}$ ($\mathcal{E}\neq\mathcal{E}^{\prime}$) being the (angular average of the) dipole-matrix-element-weighted off-diagonal elements. Here, in analogy to the momentum-space expression, $\eta_{\mathcal{E}\mathcal{E}^{\prime}}=(\eta_{\mathcal{E}}+\eta_{\mathcal{E}^{\prime}})/2$ is the decoherence rate for $\mathcal{E}\neq\mathcal{E}^{\prime}$. By Matthiessen’s rule [53], $\eta_{\mathcal{E}}=\eta_{\mathcal{E},\textrm{e-e}}+\eta_{\mathcal{E},\textrm{e-ph}}+\eta_{\mathcal{E},\textrm{e-imp}}$, where $\eta_{\mathcal{E},\textrm{e-e}}$, $\eta_{\mathcal{E},\textrm{e-ph}}$ and $\eta_{\mathcal{E},\textrm{e-imp}}$ represent the collision rates associated with the e-e, e-ph and e-imp interactions, respectively (see details in Refs. [12, 13]; they are obtained via functional derivation of their corresponding interaction terms with respect to $f(\mathcal{E})$ in energy space). $|D(k,k^{\prime},L)|^{2}$ is the angular average of the squared dipole matrix element $|d_{{\bf k}{\bf k}^{\prime}}|^{2}$ (6), i.e.,
$|D(k,k^{\prime},L)|^{2}\equiv\dfrac{1}{(4\pi)^{2}}\int d\Omega_{{\bf k}}d\Omega_{{\bf k}^{\prime}}|d_{{\bf k}{\bf k}^{\prime}}|^{2}\\ =\dfrac{2e^{2}(2\pi L)^{2}}{(kL)(k^{\prime}L)(k^{2}-k^{\prime 2})^{2}LV}\Bigg\{\dfrac{((kL)^{2}-(k^{\prime}L)^{2})^{2}-16}{4}\ln\left(\dfrac{4+|(kL)^{2}-(k^{\prime}L)^{2}|}{4+(kL-k^{\prime}L)^{2}}\right)\\ -\left[((kL)^{2}-(k^{\prime}L)^{2})^{2}+1\right]\ln\left(\dfrac{1+|(kL)^{2}-(k^{\prime}L)^{2}|}{1+(kL-k^{\prime}L)^{2}}\right)\\ -\dfrac{3}{4}\left[((kL)-(k^{\prime}L))^{2}\right]^{2}\ln\left|\dfrac{k-k^{\prime}}{k+k^{\prime}}\right|\Bigg\}.$ (16)
@@
where $k=|{\bf k}|$ and $k^{\prime}=|{\bf k}^{\prime}|$ are the magnitudes of the electron momentum, and $L$ is the system size (such as particle size or thin film thickness) in the direction of the electric field $E$. $D(k(\mathcal{E}),k(\mathcal{E}^{\prime}),L)$ has units of [C m], and thus, can be viewed as the (absolute value of the) dipole matrix element in energy space. ???. While the use of a one-dimensional infinitely deep potential well to evaluate the dipole matrix elements is a simplified model, it provides a reasonable approximation for NPs larger than a few nanometers. Indeed, in such systems, the energy spacing between adjacent states is typically $\lesssim$ 0.02-0.05 eV, comparable with the level width ($\eta$) and much smaller than the conduction bandwidth ($\gtrsim 4$ eV). This ensures a large number of accessible energy levels, justifying the approximation of discrete levels by a quasi-continuum and rendering the dipole matrix elements effectively insensitive to particle shape. Moreover, the angular averaging in Eq. (16) further mitigates the limitations of this approximation and captures the essential qualitative features of the matrix elements. Furthermore, the main effect of the exaggerated depth of the potential well is to exclude photoionization processes involving states above the vacuum level, which are anyhow not the focus of this study. A more accurate calculation of the transition dipole matrix elements can be performed using the approach in Ref. [23].

### v2

@@ Abstract
We develop an energy-space density matrix framework to investigate the interaction of extremely short optical pulses (ESPs) with transparent conducting oxides (TCOs). This approach captures not only electron populations, material polarization, and the permittivity, but also the the quantum coherences between states. Compared to traditional momentum-space models, the energy-space formulation offers substantial computational simplification while retaining accuracy. Building on but going beyond the scope of Ref. [1], we focus on dynamical features previously unexplored. Our formulation reveals clear signatures of quantum coherence in the net absorption dynamics and highlights the emergence of strong excited-state absorption under intense excitation. . Furthermore, we investigate the influence of pump pulse intensity on the local field’s duration, spectral broadening and shift, and phase induced by carrier dynamics, highlighting the absorptive nature of the nonlinear response. Our results provide a unified framework for understanding nonlinear light-matter interaction in dispersive, low-density electron systems driven far from equilibrium by intense broadband excitation.
@@ I Introduction
Initial results on this problem were published in [1], which focused on the characterization of the strength of the nonlinear response (in terms of the electron excitation dynamics, material polarization and permittivity). Instead, here, we focus on the formulational aspects of the problem, as well as on additional aspects of the dynamics that were not considered in [1]. Specifically, in Section II, we provide a detailed derivation of the DM formulation that underlies [1], including the transition dipole matrix element calculation in momentum space, and its transformation to the far more computationally-economic energy-space. Then, in Section III, we perform a thorough comparison of the strengths of the various terms appearing in the DM equations and analyze the dynamics of the coherence between the various electron states as well as the resulting repeated frequency doubling dynamics. Then, we identify the signature of the coherence in the (net) energy-integrated absorption dynamics and the emergence of significant excited state absorption within its time-integrated energy spectra; the latter is responsible for the super-linear nature of the absorption, hence, counteracting the effects of population inversion. We also justify the neglect of spontaneous emission (see Appendix. A).
@@ II.1 A derivation of a momentum-conserving energy-space density matrix formulation
Here, $\hat{\mathcal{H}}_{\textrm{e,0}}$ is the single-particle approximation for the Hamiltonian of the electrons in the lattice periodic potential, $\hat{\mathcal{H}}_{\textrm{ph,0}}$ is the unperturbed Hamiltonian of the phonon subsystem, and $\mathcal{H}_{\textrm{photon-e}}$ is the electric dipole interaction Hamiltonian
$\hat{\mathcal{H}}_{\textrm{photon-e}}=-{}{\hat{d}_{z}}\cdot E(t),$ (2)
where $E(t)$ is the local electric field, determined self-consistently from the density-matrix elements (see Section II.3). Without loss of generality, the electric field is assumed to be in the $z$-direction. $\hat{d}_{z}$ is the electric transition dipole operator along the direction of the electric field. In that regard, we neglect electron- and phonon-assisted absorption events as well as interband transitions in favor of Landau-damping-based absorption (see discussion of Ref. [26] on the analysis of Ref. [27]). Additionally, we neglect the contribution of spontaneous emission (photoluminescence), see justification in Appendix. A. $\hat{\mathcal{H}}_{\textrm{e-e}}$ is the electron-electron interaction via the Coulomb potential, $\hat{\mathcal{H}}_{\textrm{e-ph}}$ is the electron-phonon interaction via the deformation potential and $\hat{\mathcal{H}}_{\textrm{e-imp}}$ is the electron-impurity interaction.
@@
$\dot{\rho}(t)=-\dfrac{i}{\hbar}\left[\hat{\mathcal{H}}_{e},\rho\right]+\mathcal{R}(\rho)=-\dfrac{i}{\hbar}\left[\hat{\mathcal{H}}_{0}+\hat{\mathcal{H}}_{\textrm{photon-e}},\rho\right]+\mathcal{R}(\rho),$ (4)
@@
$\begin{aligned} \dot{f}_{{\bf k}}&=\mathcal{R}_{{\bf k}{\bf k}}(\rho(t))+\dfrac{i}{\hbar}E(t)\sum_{{\bf k}^{\prime}}\left[{}{(d_{z})_{{\bf k}^{\prime}{\bf k}}}\rho_{{\bf k}{\bf k}^{\prime}}-{}{(d_{z})_{{\bf k}{\bf k}^{\prime}}}\rho_{{\bf k}^{\prime}{\bf k}}\right],\end{aligned}$ (5a)
$\begin{aligned} \dot{\rho}_{{\bf k}{\bf k}^{\prime}}=\mathcal{R}_{{\bf k}{\bf k}^{\prime}}(\rho(t))-\dfrac{i}{\hbar}\left(\mathcal{E}_{{\bf k}}-\mathcal{E}_{{\bf k}^{\prime}}\right)\rho_{{\bf k}{\bf k}^{\prime}}-\dfrac{i}{\hbar}E(t){}{(d_{z})_{{\bf k}{\bf k}^{\prime}}}\left(f_{{\bf k}}(t)-f_{{\bf k}^{\prime}}(t)\right)\\ +\dfrac{i}{\hbar}E(t)\sum_{{\bf k}^{\prime\prime}\neq{\bf k},{\bf k}^{\prime}}\left[{}{(d_{z})_{{\bf k}{\bf k}^{\prime\prime}}}\rho_{{\bf k}^{\prime\prime}{\bf k}^{\prime}}-\rho_{{\bf k}{\bf k}^{\prime\prime}}{}{(d_{z})_{{\bf k}^{\prime\prime}{\bf k}^{\prime}}}\right]\end{aligned}$ $\text{for }{\bf k}\neq{\bf k}^{\prime},$ (5b)
where $\omega_{{\bf k}{\bf k}^{\prime}}\equiv\left(\mathcal{E}_{{\bf k}}-\mathcal{E}_{{\bf k}^{\prime}}\right)/\hbar$ with $\mathcal{E}_{{\bf k}}$ and $\mathcal{E}_{{\bf k}^{\prime}}$ being the renormalized energy of individual electrons and $(d_{z})_{{\bf k}{\bf k}^{\prime}}$ are the matrix elements of the electric dipole transition operator along the direction of the electric field.
We determine the momentum space dipole transition matrix elements $(d_{z})_{{\bf k}{\bf k}^{\prime}}$ from the matrix elements of the momentum operator $(p_{z})_{{\bf k}{\bf k}^{\prime}}$ and $(p^{2})_{{\bf k}{\bf k}^{\prime}}=(p_{x}^{2}+p_{y}^{2}+p_{z}^{2})_{{\bf k}{\bf k}^{\prime}}$ using the relation
${}{(d_{z})_{{\bf k}{\bf k}^{\prime}}}=-\dfrac{2ie\hbar}{(p^{2})_{{\bf k}{\bf k}}-(p^{2})_{{\bf k}^{\prime}{\bf k}^{\prime}}}{}{(p_{z})_{{\bf k}{\bf k}^{\prime}}},$ (6)
where $e$ is the electron charge. Eq. (6) is derived from the commutation relation between the position operator and the operator $\hat{{\bf p}}^{2}$ (see [35, Ch. 5]), while the matrix elements of the momentum operator $\left(p_{x,y,z}\right)_{{\bf k}{\bf k}^{\prime}}$ are determined by assuming an infinitely deep 1D potential well representing the sample geometry, as in popular earlier works [36, 37].
@@
[…]
@@
where $h_{{\bf k}{\bf k}^{\prime}}$ is a unitless phase factor such that $h_{{\bf k}{\bf k}^{\prime}}(d_{z})_{{\bf k}^{\prime}{\bf k}}=(d_{z})_{{\bf k}{\bf k}^{\prime}}$  [^4] In general, $(d_{z})_{{\bf k}^{\prime}{\bf k}}$ and $(d_{z})_{{\bf k}{\bf k}^{\prime}}$ differ by a phase factor. For our choice of a 1D symmetric potential, we have $(d_{z})_{{\bf k}^{\prime}{\bf k}}=(d_{z})_{{\bf k}{\bf k}^{\prime}}$..
@@
$\dot{\rho}_{{\bf k}{\bf k}^{\prime}}=\mathcal{R}_{{\bf k}{\bf k}^{\prime}}(\rho(t))-\dfrac{i}{\hbar}\left(\mathcal{E}_{{\bf k}}-\mathcal{E}_{{\bf k}^{\prime}}\right)\rho_{{\bf k}{\bf k}^{\prime}}-\dfrac{i}{\hbar}E(t){}{(d_{z})_{{\bf k}{\bf k}^{\prime}}}\left(f_{{\bf k}}(t)-f_{{\bf k}^{\prime}}(t)\right).$ (9)
@@
$\begin{aligned} \dfrac{\partial f_{{\bf k}}(t)}{\partial t}&=\left(\dfrac{\partial f_{{\bf k}}}{\partial t}\right)_{\textrm{e-e}}+\left(\dfrac{\partial f_{{\bf k}}}{\partial t}\right)_{\textrm{e-ph}}+\left(\dfrac{\partial f_{{\bf k}}}{\partial t}\right)_{\textrm{e-imp}}\\ &\qquad\qquad\qquad-\dfrac{i}{\hbar}E(t)\sum_{{\bf k}^{\prime}}\left[{}{(d_{z})_{{\bf k}^{\prime}{\bf k}}}\rho_{{\bf k}{\bf k}^{\prime}}(t)-{}{(d_{z})_{{\bf k}{\bf k}^{\prime}}}\rho_{{\bf k}^{\prime}{\bf k}}(t)\right]\end{aligned}$ (10a)
$\begin{aligned} \dfrac{\partial\rho_{{\bf k}{\bf k}^{\prime}}(t)}{\partial t}=-\eta_{{\bf k}{\bf k}^{\prime}}\left(\rho_{{\bf k}{\bf k}^{\prime}}-h_{{\bf k}{\bf k}^{\prime}}\rho_{{\bf k}^{\prime}{\bf k}}\right)-\dfrac{i}{\hbar}\left(\mathcal{E}_{{\bf k}}-\mathcal{E}_{{\bf k}^{\prime}}\right)\rho_{{\bf k}{\bf k}^{\prime}}-\dfrac{i}{\hbar}E(t){}{(d_{z})_{{\bf k}{\bf k}^{\prime}}}\left[f_{{\bf k}}(t)-f_{{\bf k}^{\prime}}(t)\right],\end{aligned}$ (10b)
The polarization density along the direction of the electric field can be obtained by taking the trace of the matrix multiplication of the density matrix and the dipole transition matrix, namely,
$P(t)=\dfrac{2}{V}\sum_{{\bf k},{\bf k}^{\prime}}{}{(d_{z})_{{\bf k}^{\prime}{\bf k}}}\rho_{{\bf k}{\bf k}^{\prime}}(t).$ (11)
The initial condition for the resulting system of equation includes diagonal terms at room temperature thermal distribution for the electrons (and for phonons as well), and vanishing off-diagonal terms.
@@ II.2 From $k$ to $\mathcal{E}$ space formulation
and multiply Eq. (10b) by $(d_{z})_{{\bf k}^{\prime}{\bf k}}$ and integrate over the solid angles of $\bf k$} and of $\bf k$$^{\prime}$  [^6] Note that averaging over the solid angles of $\bf k$} and $\bf k$$^{\prime}$ is fundamentally different from stochastic averaging that destroys quantum coherence. The angular integration merely accounts for isotropic contributions in momentum space and does not suppress phase coherence. Specifically, multiplying Eq. (10b) by $(d_{z})_{{\bf k}^{\prime}{\bf k}}$ yields the equation of motion for $(d_{z})_{{\bf k}^{\prime}{\bf k}}\rho_{{\bf k}{\bf k}^{\prime}}$, i.e., the $({\bf k},{\bf k}^{\prime})$ component of the polarization density. Accordingly, integration over the solid angles of $\bf k$} and $\bf k$$^{\prime}$ evaluates the total contribution to the polarization density from transitions between states with momenta $|\hbar{\bf k}|$ and $|\hbar{\bf k}^{\prime}|$. Since the source term of the equation of motion for $(d_{z})_{{\bf k}^{\prime}{\bf k}}\rho_{{\bf k}{\bf k}^{\prime}}$ is proportional to $E(t)|(d_{z})_{{\bf k}{\bf k}^{\prime}}|^{2}(f_{{\bf k}}-f_{{\bf k}^{\prime}})=E(t)|(d_{z})_{{\bf k}{\bf k}^{\prime}}|^{2}\big[f(\mathcal{E}(|{\bf k}|))-f(\mathcal{E}(|{\bf k}^{\prime}|))\big]$, which has a definite sign (in phase), $\int d\Omega_{{\bf k}}d\Omega_{{\bf k}^{\prime}}d_{{\bf k}^{\prime}{\bf k}}\rho_{{\bf k}{\bf k}^{\prime}}$ produces a finite contribution to the polarization density. (see details in Ref. [28]). The photon-e interaction term, the off-diagonal density matrix equation of motion Eq. (10b), and the polarization density Eq. (11) thus, respectively, become
@@
with $\overline{d_{\mathcal{E}^{\prime}\mathcal{E}}\rho_{\mathcal{E}\mathcal{E}^{\prime}}}\equiv\dfrac{1}{(4\pi)^{2}}\displaystyle\int d\Omega_{{\bf k}}d\Omega_{{\bf k}^{\prime}}{}{(d_{z})_{{\bf k}^{\prime}{\bf k}}}\rho_{{\bf k}{\bf k}^{\prime}}$ ($\mathcal{E}\neq\mathcal{E}^{\prime}$) being the (angular average of the) dipole-matrix-element-weighted off-diagonal elements. Here, in analogy to the momentum-space expression, $\eta_{\mathcal{E}\mathcal{E}^{\prime}}=(\eta_{\mathcal{E}}+\eta_{\mathcal{E}^{\prime}})/2$ is the decoherence rate for $\mathcal{E}\neq\mathcal{E}^{\prime}$. By Matthiessen’s rule [56], $\eta_{\mathcal{E}}=\eta_{\mathcal{E},\textrm{e-e}}+\eta_{\mathcal{E},\textrm{e-ph}}+\eta_{\mathcal{E},\textrm{e-imp}}$, where $\eta_{\mathcal{E},\textrm{e-e}}$, $\eta_{\mathcal{E},\textrm{e-ph}}$ and $\eta_{\mathcal{E},\textrm{e-imp}}$ represent the collision rates associated with the e-e, e-ph and e-imp interactions, respectively ( [^7] Here, the e-e interaction in Eq. (10a) includes only the normal processes when determining the electron distribution. Upon angular averaging, the contribution from these normal scattering processes vanishes because of overall momentum conservation. Instead, Umklapp processes contribute to the relaxation of the polarization (or equivalently, the induced current). For ITO, this ratio is presently unknown; however, it is likely to be of order unity, $\mathcal{O}(1)$ [83]., see also details in Refs. [12, 13]; they are obtained via functional derivation of their corresponding interaction terms with respect to $f(\mathcal{E})$ in energy space  [^8] Note that the expression in [13], namely, Eq. (B6), had a small typo; the expression in the rectangular brackets on the 2nd line should have read $\left(1+\frac{q^{2}_{TF}}{2k^{2}}\right)^{2}-1$.). $|D(k,k^{\prime},L)|^{2}$ is the angular average of the squared dipole matrix element $|(d_{z})_{{\bf k}{\bf k}^{\prime}}|^{2}$ (6), i.e.,
$|D(k,k^{\prime},L)|^{2}\equiv\dfrac{1}{(4\pi)^{2}}\int d\Omega_{{\bf k}}d\Omega_{{\bf k}^{\prime}}|{}{(d_{z})_{{\bf k}{\bf k}^{\prime}}}|^{2}\\ =\dfrac{2e^{2}(2\pi L)^{2}}{(kL)(k^{\prime}L)(k^{2}-k^{\prime 2})^{2}LV}\Bigg\{\dfrac{((kL)^{2}-(k^{\prime}L)^{2})^{2}-16}{4}\ln\left(\dfrac{4+|(kL)^{2}-(k^{\prime}L)^{2}|}{4+(kL-k^{\prime}L)^{2}}\right)\\ -\left[((kL)^{2}-(k^{\prime}L)^{2})^{2}+1\right]\ln\left(\dfrac{1+|(kL)^{2}-(k^{\prime}L)^{2}|}{1+(kL-k^{\prime}L)^{2}}\right)\\ -\dfrac{3}{4}\left[((kL)-(k^{\prime}L))^{2}\right]^{2}\ln\left|\dfrac{k-k^{\prime}}{k+k^{\prime}}\right|\Bigg\}.$ (16)
@@
where $k=|{\bf k}|$ and $k^{\prime}=|{\bf k}^{\prime}|$ are the magnitudes of the electron momentum, and $L$ is the system size (such as particle size or thin film thickness) in the direction of the electric field $E$. $D(k(\mathcal{E}),k(\mathcal{E}^{\prime}),L)$ has units of [C m], and thus, can be viewed as the (absolute value of the) dipole matrix element in energy space. While the use of a one-dimensional infinitely deep potential well to evaluate the dipole matrix elements is a simplified model, it provides a reasonable approximation for NPs larger than a few nanometers. Indeed, in such systems, the energy spacing between adjacent states is typically $\lesssim 0.02-0.05$ eV, comparable with the level width ($\eta$) and much smaller than the conduction bandwidth ($\gtrsim 4$ eV). This ensures a large number of accessible energy levels, justifying the approximation of discrete levels by a quasi-continuum and rendering the dipole matrix elements effectively insensitive to particle shape. Moreover, the angular averaging in Eq. (16) further mitigates the limitations of this approximation and captures the essential qualitative features of the matrix elements. Furthermore, the main effect of the exaggerated depth of the potential well is to exclude photoionization processes involving states above the vacuum level, which are anyhow not the focus of this study. A more accurate calculation of the transition dipole matrix elements can be performed using the approach in Ref. [23].